OpenStudy (anonymous):

How far does an object travel horizontally at 35ft/s use the equation x(t)=vcos(45)t

OpenStudy (anonymous):

x(t)=vcos(35)t

OpenStudy (anonymous):

I think they're saying \[ v =35 \\ x(t) = 35\cos(45^\circ )t \]

OpenStudy (anonymous):

Yes i cant find a way to do it! Please help

OpenStudy (anonymous):

You need to figure out \(t\). I assume there's gravity involved.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[ y(t) = \frac{1}{2}at^2+v_{y0}t+y_0 \]

OpenStudy (anonymous):

\[ y_0 = 0 \\ v_{y0} = \sin(45^\circ) \\ a = -9.8 \]

OpenStudy (anonymous):

It's a quadratic equation. Find the roots. One of them is obviously \(0\)

OpenStudy (anonymous):

\[R = \frac{ u^2 \sin2\theta }{ g }\] u = 35 theta = 45 2theta = 90

OpenStudy (anonymous):

\[R = \frac{ u^2 *1}{ g }\]

OpenStudy (anonymous):

Whoops, my \(a\) was in metric... you should use \(a=32\)

OpenStudy (anonymous):

Oh ok that confused me

OpenStudy (anonymous):

\[ y(t)= \left[ \frac{1}{2}(32)t-35\sin(45^\circ) \right]t \]

OpenStudy (anonymous):

@El_Pollo Can you solve for \(t\)?

OpenStudy (anonymous):

Set \(y(t) = 0\)

OpenStudy (anonymous):

t=1.54

OpenStudy (anonymous):

Now find \(x(1.54)\) given the equation above.